H2C2O4 + 2Cr3+2CO2 + 2Cr2++ 2H+


In the above redox reaction, use oxidation numbers to identify the element oxidized, the element reduced, the oxidizing agent and the reducing agent.
name of the element oxidized:
name of the element reduced:
formula of the oxidizing agent:
formula of the reducing agent:


2F- + HgO+ H2OHg + F2+ 2OH-

In the above redox reaction, use oxidation numbers to identify the element oxidized, the element reduced, the oxidizing agent and the reducing agent.

name of the element oxidized:
name of the element reduced:
formula of the oxidizing agent:
formula of the reducing agent:



Si + 2F2+ 6OH-4F- + SiO32-+ 3H2O
In the above redox reaction, use oxidation numbers to identify the element oxidized, the element reduced, the oxidizing agent and the reducing agent.

name of the element oxidized: silicon
name of the element reduced:
formula of the oxidizing agent:
formula of the reducing agent:
1
Expert's answer
2020-01-20T05:11:09-0500

Si + 2F2 + 6OH- = 4F- + SiO32- + 3H2O

Si - 4e = Si4+

F2 + 2e = 2F-

name of the element oxidized: silicon Si

name of the element reduced: fluorine F

formula of the oxidizing agent: F2

formula of the reducing agent: Si


2F- + HgO + H2O = Hg + F2 + 2OH-

2F- -2e = F2

Hg2+ + 2e = Hg

name of the element oxidized: fluorine F

name of the element reduced: mercury Hg

formula of the oxidizing agent: HgO

formula of the reducing agent: F-


H2C2O4 + 2Cr3+ = 2CO2 + 2Cr2+ + 2H+

2C3+ - 2e = 2C4+

Cr3+ + 1e = Cr2+

name of the element oxidized: carbon C

name of the element reduced: chromium Cr

formula of the oxidizing agent: Cr3+

formula of the reducing agent: H2C2O4